# A Benchmark for End-to-End Zero-Shot Biomedical Relation Extraction with LLMs: Experiments with OpenAI Models

### v1

@@ 3.3 Evaluation
As a task, RE is complex, and there are many reasonable ways to measure performance. This has led to a proliferation of measures, but also confusion and conflation of them — so much so that rigorous study of the issue has been made [25]. Unfortunately, the state of affairs has only worsened as (1) researchers have not heeded this work, (2) papers have faded descriptions of details necessary for reproducibility, (3) EL RE has been introduced, and (4) seq2seq methods, which have grown in popularity, lend themselves to new performance measures.
Among the three main methods that have been published for EL RE — JEREX [26], seq2rel [19], and BioGPT [8] — no two calculate F1 in the same way. JEREX measures performance very strictly: a predicted relation is considered correct if it matches a gold relation exactly, and entities within the relations are judged correct when mentioned boundaries are correct. Seq2rel’s “strict” measure is similar, except that rather than entity mentions being judged on boundary correctness, they are judged on whether the predicted strings match gold entity mention strings, and duplicate gold mention strings are collapsed to a single mention. JEREX correctness therefore implies seq2rel “strict” correctness, but not vice versa. We note that since seq2rel uses a copy mechanism that points directly to tokens, nothing prevents them from making an exact comparison with JEREX. However, Giorgi et al. [19] compared their performance with JEREX using slightly different “strict” metrics as indicated earlier. They additionally use a “relaxed” measure of correctness that only requires a majority of predicted entity mentions to match that of a gold entity.
The generative approach of BioGPT lends itself to the extraction of a single entity mention rather than all of them, and therefore Luo et al. [8] deem a predicted relation correct if the extracted mentions match the longest mention in the dataset, rather than the example text. Further distinguishing their performance measure from previous papers, Luo et al. [8] filter examples with no gold relations from the dataset. Despite these differences, they compare their performance with seq2rel. At this point, it is not clear, on any dataset, which of these methods has the highest performance; nor is it clear that they all can be compared with one another, even if done with utmost care. To make matters worse, Jahan et al.[21] do not describe their evaluation methodology or provide code.
@@ 3.4 Datasets
Table 1 describes the basic properties of the datasets we studied. Three of them contain EL relations; these datasets naturally contain relations with entity mentions across multiple sentences. The remaining four datasets contain intra-sentence ML relations, though relation types may be more easily extracted when the surrounding context is available.
@@
The ADE dataset [27] consists of sentences extracted from MEDLINE case reports describing adverse effects resulting from drug use, extracted from medical case reports. It contains 2 entity types: drugs and adverse effects and 1 relation type, adverse drug event. There is no official split of the dataset.
@@
DCE [1] documents the efficacy of drug combination therapies, presenting a unique RE challenge in that relations contain a variable number of entity types. Each instance consists of an abstract, within which a focal sentence is identified that contains multiple drug references. The drug references are classified as either being positive, for a beneficial drug combination, non-positive, for a combination with a neutral or negative effect, or non-combination, when the drugs are not given in combination. Following the practice of the original authors, DCE performance is evaluated using two metrics: Positive Combination F1 score and Any Combination F1 score. The Positive Combination F1 treats the relation type positive as the positive class, while the Any Combination F1 score lumps positive and non-positive relation types together, and treats them as the positive class.
@@
The primary aim of ChemProt [28] is to extract intra-sentence relations between chemical compounds and proteins/genes from biomedical abstracts. Relation types holding between these entities can be described as upregulator, downregulator, agonist, antagonist, or substrate of. Over 25% of abstracts contain no relations.
Datasets Type Input Type ##\## Entity Types ##\## Predicates Examples w/o relations
ADE [27] Mention-Level Sentence 2 1 No
DCE [1] Abstract 1 3 Yes
ChemProt [28] Abstract 2 5 Yes
DDI [29] Abstract 4 4 Yes
CDR [3] Entity-Level Abstract 2 1 No
GDA [30] Abstract 2 1 No
BioRED [2] Abstract 4 8 Yes
@@
DDI [29] annotates intra-sentence interactions between four types of pharmacological substances: brand-name drugs, generic drugs, drug categories, and substances not approved for human use. Drug-drug interactions are either descriptions of pharmacokinetic mechanisms, descriptions of effect/pharmacodynamic mechanisms, recommendations about drug combinations, or documented interactions without additional details. Nearly two-thirds of instances contain no relations. The CDR [3] and GDA [30] datasets respectively annotate diseases induced by chemicals/drugs or associated with genes in PubMed abstracts. Both datasets contain EL relations with a single relation type.
@@
BioRED [2] dataset annotates eight non-directional relation types holding between genes, gene variants, chemicals, and diseases. The relation types are positive correlation, negative correlation, association, binding, co-treatment, drug interaction, comparison, and conversion; certain relation types are only valid for a subset of all combinations of entity types. Instances in BioRED often contain many relations, sometimes in excess of 90. Presumably, for this reason, there are only 100 test instances.
@@
We show ZS prompts for ChemProt and CDR in Table 4. All prompts for seven biomedical datasets are in https://github.com/bionlproc/ZeroShotRE/tree/main/prompts.
4 RESULTS AND DISCUSSION
@@
We present our results for GPT-4 and o1 shown in Table 2. Our method varied substantially in performance across datasets, with it being much higher in ADE, DCE, CDR, and GDA than ChemProt, DDI, and BioRED. Separating these two groups of datasets is the number of relations — high performance was achieved on datasets with 1-2 relation types, while low performance was obtained on datasets with 4-8 relation types. The low-performance datasets additionally tended to have higher numbers of entity types. These may be spurious correlations — future work could test this hypothesis by slicing these datasets by entity/relation types and checking whether performance improves.
We discussed in Section 3.3 that a valid comparison of performance between current biomedical RE methods is generally not possible. However, all of the performance measures are obviously positively correlated, so we collate performance from other publications with reasonably transparent evaluation methodology in Table 3. We find that supervised methods using far smaller LMs perform similarly or better than our method, an unsurprising result [10]. However, due to the aforementioned difficulties of comparison, the only obvious discrepancy occurs in ChemProt, where our method fared poorly.
Datasets GPT-4 (Inferred) GPT-4 (Explicit) OpenAI o1 (Inferred)
$P$ $R$ $F_{1}$ $P$ $R$ $F_{1}$ $P$ $R$ $F_{1}$
ADE [27] 75.3 60.4 67.0 76.7 62.8 69.1 73.5 62.8 67.7
DCE (Positive combination) [1] 58.9 68.7 63.4 61.3 66.7 63.9 61.5 74.7 67.5
DCE (Any combination) [1] 55.6 76.1 64.2 49.2 71.8 58.4 69.6 74.6 72.1
ChemProt [28] 24.1 24.6 24.3 19.7 23.0 21.2 37.0 20.7 26.5
DDI [29] 27.7 33.6 30.4 27.6 33.2 30.1 46.1 29.3 35.8
CDR [3] 48.9 42.3 45.3 49.3 42.4 45.6 46.6 41.2 43.7
GDA [30] 46.0 63.4 53.3 46.1 65.2 54.0 40.2 57.6 47.3
BioRED [2] 12.6 7.1 9.1 15.1 7.3 9.8 30.8 18.6 23.2
Average 41.7 43.4 41.9 41.4 43.3 41.6 48.5 43.6 44.9
@@
Table 2: Main results for RE experiments. As DCE is evaluated in two ways (see 3.4), their performance values are averaged before being included in the calculation for the "Average" row. Since the GPT-4 Inferred and Explicit methods differ only slightly, we chose to apply only the Inferred method to OpenAI o1. As a result, OpenAI o1 surpasses GPT-4 in most datasets.
We conducted a detailed error analysis on all datasets to glean insights into the pitfalls of ZSRE. We first note which aspects of RE were highly successful. GPT-4 was nearly perfectly faithful to the structured schema we described in our templates and generated entity and relation types were nearly always selected from the set of types we described in the templates. Predicted entity mentions were usually assigned the correct entity type as well. Predicted mentions are rarely not present in the text. Last, it was uncommon for relation types to be incorrect when entities participating in relations were correct.
Compared to the o1 model, GPT-4 generally underperforms except on the ADE, CDR and GDA datasets—all of which are simpler, containing only a single relation type. One possible explanation is the explicit accommodation in o1 for tasks with more complexities with reasoning ability. Consequently, o1 achieves better performance than GPT-4 on the ChemProt, BioRED, and DDI datasets, each of which contains 4–8 distinct relation types. On average, the o1 model is over 3 F1 points better than GPT-4 and the biggest relative change is seen in BioRED where o1 more than doubles the GPT-4 performance. The precision-recall patterns are identical for GPT-4 and o1 except for ChemProt and DDI where precision is better than recall for o1, while GPT-4 has more recall than precision.
Both models generally under-predict relations when an instance contains more than a few gold relations. Fig 2 depicts this pattern for CDR on GPT-4, a representative example. It shows that the average number of relations predicted per test instance lags further behind the number of gold relations as the number of gold relations increases, and that this naturally results in decreased recall. We attribute this to the fact that generative models tend to perform worse with long sequences [31, 32].
A common error across datasets and models we encountered was that of partial matching entity mentions, in which a predicted relation nearly matches a gold relation, but predicted mentions either include extra words not found in gold mentions or exclude words found in them. For example, we extracted the incorrect chemical-disease relation (chemical: methamphetamine, disease: methamphetamine-induced psychosis), which would have been correct had we extracted the disease as psychosis. Future research should focus on extracting correct boundaries for entity mentions, as this was a large source of error.
Datasets ADE DCE ChemProt DDI CDR GDA BioRED
Yan et al. [33] 83.2 - - - - - -
@@
BioGPT [8] - - - 40.8 46.2 - -
PURE [35] - - 69.0[36] - - - -
@@
Table 3: Comparison of performance (F1) of OpenAI zero-shot scores with previous finetuned methods. Note that F1 does not have identical meaning across methods (see Section 3.3). We do not include any performance numbers from papers that do not describe the evaluation methodology clearly. Superscripts P and A refer to the “Positive” and “Any Combination” evaluation settings, as described in Section 3.4. Superscripts $\dagger$ and $\ddagger$ refer to the “strict” and “relaxed” evaluations described in Section 3.4.
@@
Frequently, false positives appear to be correct relations missed by the annotators. This has been previously documented in RE datasets [37, 38] and has been shown to artificially severely deflate performance. Given the high frequency of missed relations, it may be prudent to re-annotate biomedical benchmark RE datasets in the mode of Tan et al. [38]
@@
For the most part, the two models predicted mentions that either were exact spans of source text or concatenated discontinuous spans. However, in some cases, they used domain knowledge, predicting a text string not found in the source text. In one extracted relation from GDA, we predicted the entity interleukin-10, which did not appear in the text in this form, whereas the gold version was interleukin (IL)-10.
@@
In addition to patterns of errors holding across datasets, errors arose particular to specific datasets. Our method largely failed on BioRED, with frequent, obviously incorrect, predicted entity mentions and types as well as relations claiming opposite relationships between two given entities. In DCE, drug combinations were frequently missing drugs that participate in a relation. Also, relation types were often incorrectly assigned; we suspect that this is caused by the domain-specific knowledge, like the interpretation of lab result quantities, sometimes required to correctly assign relation type.
All results for the use of OpenAI GPT models for publicly available datasets since GPT-3 [39] come with a caveat: we do not know what data the models were trained on. These models are trained using massive amounts of scraped web data, and all datasets we used are available on the web. It is possible that ChatGPT models was trained on these datasets, which would invalidate the performance values we obtained. However the very low scores obtained for BioRED and ChemProt indicate that this contamination may not have happened. The development of methods that can deduce whether a LM has been trained on a specific dataset would be invaluable.
5 CONCLUSION
The OpenAI LLMs, including the advanced reasoning model o1, have demonstrated ZS capabilities similar to smaller, finetuned models on simple tasks. The high-value, complex task of biomedical RE presents a more significant challenge in the ZS arena. We demonstrate that across diverse datasets, ZS performance of LLMs can be competitive, and highlight areas for improvement. We also demonstrate that LLMs equipped with reasoning capabilities can achieve better performances. Our approach trades effort and expertise in dataset curation and NLP modeling, which may increase access to the medical community. It also shows that for cases where the density of relations per instance is high, LLMs have a much tougher time and have a long way to go for end-to-end ZSRE.

### v3

@@ 2.3 Evaluation
As a task, RE is complex, and there are many reasonable ways to measure performance. This has led to a proliferation of measures, but also confusion and conflation of them — so much so that rigorous study of the issue has been made Taillé et al. (2020). Unfortunately, the state of affairs has only worsened as (1) researchers have not heeded this work, (2) papers have faded descriptions of details necessary for reproducibility, (3) EL RE has been introduced, and (4) seq2seq methods, which have grown in popularity, lend themselves to new performance measures.
Among the three main methods that have been published for EL RE — JEREX Eberts and Ulges (2021), seq2rel Giorgi et al. (2022), and BioGPT Luo et al. (2022b) — no two calculate F1 in the same way. JEREX measures performance very strictly: a predicted relation is considered correct if it matches a gold relation exactly, and entities within the relations are judged correct when mentioned boundaries are correct. Seq2rel’s “strict” measure is similar, except that rather than entity mentions being judged on boundary correctness, they are judged on whether the predicted strings match gold entity mention strings, and duplicate gold mention strings are collapsed to a single mention. JEREX correctness therefore implies seq2rel “strict” correctness, but not vice versa. We note that since seq2rel uses a copy mechanism that points directly to tokens, nothing prevents them from making an exact comparison with JEREX. However, Giorgi et al. (2022) compared their performance with JEREX using slightly different “strict” metrics as indicated earlier. They additionally use a “relaxed” measure of correctness that only requires a majority of predicted entity mentions to match that of a gold entity.
The generative approach of BioGPT lends itself to the extraction of a single entity mention rather than all of them, and therefore Luo et al. (2022b) deem a predicted relation correct if the extracted mentions match the longest mention in the dataset, rather than the example text. Further distinguishing their performance measure from previous papers, Luo et al. (2022b) filter examples with no gold relations from the dataset. Despite these differences, they compare their performance with seq2rel. At this point, it is not clear, on any dataset, which of these methods has the highest performance; nor is it clear that they all can be compared with one another, even if done with utmost care. To make matters worse, Jahan et al. (2023) do not describe their evaluation methodology or provide code.
Datasets Type Input Type #\# Entity Types #\# Predicates Examples w/o relations
ADE Gurulingappa et al. (2012) Mention-Level Sentence 2 1 No
DCE Tiktinsky et al. (2022) Abstract 1 3 Yes
ChemProt Krallinger et al. (2017) Abstract 2 5 Yes
DDI Herrero-Zazo et al. (2013) Abstract 4 4 Yes
CDR Li et al. (2016) Entity-Level Abstract 2 1 No
GDA Wu et al. (2019) Abstract 2 1 No
BioRED Luo et al. (2022a) Abstract 4 8 Yes
@@ 2.4 Datasets
Table 1 shows the basic properties of the datasets we studied. Three of them contain EL relations; these datasets naturally contain relations with entity mentions across multiple sentences. The remaining four datasets contain intra-sentence ML relations, though relation types may be more easily extracted when the surrounding context is available.
@@
The ADE dataset Gurulingappa et al. (2012) consists of sentences extracted from MEDLINE case reports describing adverse effects resulting from drug use, extracted from medical case reports. It contains two entity types: drugs and adverse effects and one relation type, adverse drug event. There is no official split of the dataset.
@@
DCE Tiktinsky et al. (2022) documents the efficacy of drug combination therapies, presenting a unique RE challenge in that relations contain a variable number of entity types. Each instance consists of an abstract, within which a focal sentence is identified that contains multiple drug references. The drug references are classified as either being positive, for a beneficial drug combination, non-positive, for a combination with a neutral or negative effect, or non-combination, when the drugs are not given in combination. Following the practice of the original authors, DCE performance is evaluated using two metrics: Positive Combination F1 score and Any Combination F1 score. The Positive Combination F1 treats the relation type positive as the positive class, while the Any Combination F1 score lumps positive and non-positive relation types together, and treats them as the positive class.
@@
The primary aim of ChemProt Krallinger et al. (2017) is to extract intra-sentence relations between chemical compounds and proteins/genes from biomedical abstracts. Relation types holding between these entities can be described as upregulator, downregulator, agonist, antagonist, or substrate of. Over 25% of abstracts contain no relations.
@@
DDI Herrero-Zazo et al. (2013) annotates intra-sentence interactions between four types of pharmacological substances: brand-name drugs, generic drugs, drug categories, and substances not approved for human use. Drug-drug interactions are either descriptions of pharmacokinetic mechanisms, descriptions of effect/pharmacodynamic mechanisms, recommendations about drug combinations, or documented interactions without additional details. Nearly two-thirds of instances contain no relations. The CDR Li et al. (2016) and GDA Wu et al. (2019) datasets respectively annotate diseases induced by chemicals/drugs or associated with genes in PubMed abstracts. Both datasets contain EL relations with a single relation type.
@@
BioRED Luo et al. (2022a) dataset annotates eight non-directional relation types holding between genes, gene variants, chemicals, and diseases. The relation types are positive correlation, negative correlation, association, binding, co-treatment, drug interaction, comparison, and conversion; certain relation types are only valid for a subset of all combinations of entity types. Instances in BioRED often contain many relations, sometimes in excess of 90. Presumably, for this reason, there are only 100 test instances.
@@
We show ZS prompts for ChemProt and CDR in Table 4 in the Appendix. Prompts for all seven biomedical datasets are made available here: https://github.com/bionlproc/ZeroShotRE/tree/main/prompts.
3 Results
@@
We present results for GPT-4, OpenAI o1, and GPT-OSS-120B in Table 2. Performance varies substantially across datasets, with consistently higher scores on ADE, DCE, CDR, and GDA compared to ChemProt, DDI, and BioRED. A clear pattern emerges: datasets with only 1 or 2 relation types yield much higher performance, while those with 4–8 relation types, often accompanied by a larger set of entity types, show lower performance. These observations may reflect spurious correlations; future work could test this hypothesis by slicing datasets by entity and relation types to examine whether performance improves.
When comparing models, GPT-4 generally underperforms relative to both o1 and GPT-OSS-120B. GPT-OSS consistently surpasses GPT-4 across all datasets, while o1 also outperforms GPT-4 except on ADE, CDR, and GDA—datasets that are simpler and contain only a single relation type. One explanation is that o1 and GPT-OSS-120B better accommodate tasks with greater relational and entity complexity through more robust reasoning. As a result, both surpass GPT-4 on ChemProt, BioRED, and DDI, each of which contains 4–8 distinct relation types. On average, o1 is over 3 F1 points better than GPT-4, with BioRED showing the largest relative gain (more than doubling GPT-4 performance). GPT-OSS-120B achieves the strongest overall results, attaining the best F1 scores on most datasets. It is encouraging to see a open-weights model that is better than large closed models (by an average four F1 points over the o1 model). Precision–recall trends are largely consistent across models, though o1 and GPT-OSS-120B tend to favor precision on ChemProt and DDI, while GPT-4 exhibits relatively higher recall.
Datasets GPT-4 (Inferred) GPT-4 (Explicit) OpenAI o1 (Inferred) GPT-OSS-120B (Inferred)
$P$ $R$ $F_{1}$ $P$ $R$ $F_{1}$ $P$ $R$ $F_{1}$ $P$ $R$ $F_{1}$
ADE Gurulingappa et al. (2012) 75.3 60.4 67.0 76.7 62.8 69.1 73.5 62.8 67.7 76.7 68.8 72.5
DCE (Pos.) Tiktinsky et al. (2022) 58.9 68.7 63.4 61.3 66.7 63.9 61.5 74.7 67.5 64.7 73.3 68.8
DCE (Any) Tiktinsky et al. (2022) 55.6 76.1 64.2 49.2 71.8 58.4 69.6 74.6 72.1 70.6 77.0 73.7
ChemProt Krallinger et al. (2017) 24.1 24.6 24.3 19.7 23.0 21.2 37.0 20.7 26.5 28.9 31.0 30.0
DDI Herrero-Zazo et al. (2013) 27.7 33.6 30.4 27.6 33.2 30.1 46.1 29.3 35.8 36.2 53.3 43.1
CDR Li et al. (2016) 48.9 42.3 45.3 49.3 42.4 45.6 46.6 41.2 43.7 52.1 48.5 50.2
GDA Wu et al. (2019) 46.0 63.4 53.3 46.1 65.2 54.0 40.2 57.6 47.3 45.5 67.4 54.4
BioRED Luo et al. (2022a) 12.6 7.1 9.1 15.1 7.3 9.8 30.8 18.6 23.2 27.0 17.9 21.5
Average 41.7 43.4 41.9 41.4 43.3 41.6 48.5 43.6 44.9 47.7 51.7 49.0
@@
Table 2: Main results for end-to-end ZSRE experiments. As DCE is evaluated in two ways (see Section 2.4), their performance values are averaged before being included in the calculation for the “Average” row.
As discussed in Section 2.3, a valid comparison of performance between current biomedical RE methods is generally not possible. However, all of the performance measures are obviously positively correlated, so we collate performance from other publications with reasonably transparent evaluation methodology in Table 3. We find that supervised methods using far smaller LMs perform similarly or better than our method, an unsurprising result Wadhwa et al. (2023). However, due to the aforementioned difficulties of comparison, the only obvious discrepancy occurs in ChemProt, where our method fared poorly. ChemProt encodes fine-grained and mechanistic relations between chemicals and proteins where the predicates are semantically close. This may require explicit learning of subtle lexical cues and biochemical context, which is difficult in the ZS setting, leading to the large performance gap relative to the supervised score.
Methods ADE DCE ChemProt DDI CDR GDA BioRED
Yan et al. Yan et al. (2021) 83.2 - - - - - -
@@
BioGPT Luo et al. (2022b) - - - 40.8 46.2 - -
PURE Zhong and Chen (2021) - - 69.0 - - - -
@@
Table 3: Comparison of performance (F1) of OpenAI ZS scores with previous finetuned methods. Note that F1 does not have identical meaning across methods (see Section 2.3). Superscripts P and A refer to the “Positive” and “Any Combination” evaluation settings for DCE Jiang and Kavuluru (2023). Superscripts $\dagger$ and $\ddagger$ refer to the “strict” and “relaxed” evaluations described in Section 2.3.
We analyzed errors from all datasets to glean insights into the pitfalls of ZSRE. We first note which aspects of RE were highly successful. GPT-4 was nearly perfectly faithful to the structured schema we described in our templates and generated entity and relation types were nearly always selected from the set of types we described in the templates. Predicted entity mentions were usually assigned the correct entity type as well. Predicted mentions are rarely not present in the text. Last, it was uncommon for relation types to be incorrect when entities participating in relations were correct.
All models tend to under-predict relations when an instance contains more than a few gold relations. Figure 1 in the Appendix depicts this pattern for CDR on GPT-4, a representative example. It shows that the average number of relations predicted per test instance lags further behind the number of gold relations as the number of gold relations increases, and that this naturally results in decreased recall. We attribute this to the observation that generative models tend to perform worse with long sequences Hochreiter et al. (2001); Li et al. (2023a).
A common error across datasets and models we encountered was that of partial matching entity mentions, in which a predicted relation nearly matches a gold relation, but predicted mentions either include extra words not found in gold mentions or exclude words found in them. For example, we extracted the incorrect chemical-disease relation (chemical: methamphetamine, disease: methamphetamine-induced psychosis), which would have been correct had we extracted the disease as psychosis. Future research should focus on extracting correct boundaries for entity mentions, as this was a major source of error.
@@
Frequently, false positives appear to be correct relations missed by the annotators. This has been previously documented in RE datasets Tran and Kavuluru (2019); Tan et al. (2022) and has been shown to artificially deflate performance. Given the high frequency of missed relations, it may be prudent to re-annotate biomedical benchmark RE datasets in the mode of Tan et al. (2022).
@@
For the most part, models predicted mentions that either were exact spans of source text or concatenated discontinuous spans. However, in some cases, they used domain knowledge, predicting a text string not found in the source text. In one extracted relation from GDA, we predicted the entity interleukin-10, which did not appear in the text in this form, whereas the gold version was interleukin (IL)-10.
@@
Besides error patterns holding across datasets, errors arose particular to specific datasets. Our method largely failed on BioRED, with frequent, obviously incorrect, predicted entity mentions and types as well as relations claiming opposite relationships between two given entities. In DCE, drug combinations were frequently missing drugs that participate in a relation. Also, relation types were often incorrectly assigned; we suspect that this is caused by the domain-specific knowledge, like the interpretation of lab result quantities, sometimes required to correctly assign relation type.
4 Discussion
In the LLM era, a fundamental question is whether and in what settings can we simply use ZS predictions from frontier LLMs without the tedious and expensive creation of training data and custom supervised models. The answer to this is heavily dependent on particular task on hand in terms of expectations on recall and precision and the consequences of false positives/negatives. For example, if the relations are being used for knowledge discovery, focusing more on precision can minimize creation of misleading hypotheses. However, to conduct systematic reviews on information encoded in the relations (e.g., drug–side-effects), the relations extracted should have high recall.
As a first step to assess the general end-to-end ZSRE competence of LLMs, we created a new benchmark and conducted experiments with frontier LLMs. Our high level takeaway is that for shorter instances with fewer relations (less information density), ZSRE is closer to fully supervised models; this is more so if the entities are shorter (fewer tokens) such as mostly single token drug entities in DCE and ADE datasets. Entity complexity is also high in ChemProt dataset, which may have contributed to the vast discrepancy (40 points in F1) between supervised and ZS performances. Relatively, high density longer inputs (e.g., BioRED) lead to almost unusable performance at this point. Further assessments that account for partial matches of entities may be warranted but unless that is done carefully, the results may not be meaningful. For instance, partial matches that do not involve the head word of an entity phrase are mostly incorrect and misleading at best.
All results using OpenAI GPT models for publicly available datasets since GPT-3 Brown et al. (2020) come with a caveat: we do not know what data the models were trained on. These LLMs train on massive amounts of scraped web data, and most datasets we used are available on the web in some form. It is possible that they may have been trained on these datasets. However the very low scores obtained for BioRED and ChemProt indicate that this contamination is unlikely. Please note that here we are not focusing on the textual inputs (sans labels) being part of the LLM pre-training corpora; this was shown to not cause any contamination in general Li et al. (2024). The focus here is on whether particular task-specific labels were part of the training process. Recently, methods to detect membership of specific texts in the pretraining corpora of LLMs have been introduced Rastogi et al. (2025) but foolproof tools that assess whether ground truth labels of supervised tasks are leaked are not available. Recent evidence shows that label contamination is possible for question answering tasks where the answer is a single word or a short phrase, but if rich, structured outputs (as in end-to-end RE) are required, memorization is not as prevalent Wang et al. (2025). While label contamination should be kept in mind, that potential should not be grounds for not evaluating ZS performances; other teams have been exploring the same with OpenAI models Li et al. (2023b); Zhang et al. (2024), albeit not in an end-to-end manner. In light of recent calls for action Jacovi et al. (2023), we have encrypted the datasets using in this project before hosting them in our GitHub space. Others can still reproduce our results by running our scripts, which would decrypt them on the fly in a programmatic manner.
5 Limitations
Our results and associated implications apply only to English datasets, but we hope others will follow up with benchmarks in other languages. With regards to LLMs studied, we only considered OpenAI models while there are more frontier options (e.g., from Anthropic, Google, and xAI), which we could not work with due to time and cost constraints. Also, there are datasets with more complex annotation schemes than BioRED such as the cancer genetics and pathway curation datasets of the BioNLP 2013 shared task Nédellec et al. (2013). However, the download links for them are not active and additional efforts are needed to carefully recover and experiment with them. Finally, as we already acknowledged in Section 4, there is some potential risk of label contamination although we believe it is minimal for RE tasks where memorization is nontrivial.
Acknowledgment
This work is supported by the U.S. NIH National Library of Medicine through grant R01LM013240. The content is solely the responsibility of the authors and does not necessarily represent the official views of the NIH.
Appendix A Appendix
In Table 4, we give examples of ZS prompts for ChemProt and CDR datasets each describing the task, predicate label definitions, along with JSON output templates with some dummy filled in examples. In Figure 1, we show how the number of predicted relations does not keep up as the number of gold relations increases (x-axis) and hence recall decreases as the number of gold relations increase in an instance.